# Incorporating Higher-order Structural Information for Graph Clustering

### 17 Mar 2024 v1

@@ Abstract
Clustering holds profound significance in data mining. In recent years, graph convolutional network (GCN) has emerged as a powerful tool for deep clustering, integrating both graph structural information and node attributes. However, most existing methods ignore the higher-order structural information of the graph. Evidently, nodes within the same cluster can establish distant connections. Besides, recent deep clustering methods usually apply a self-supervised module to monitor the training process of their model, focusing solely on node attributes without paying attention to graph structure. In this paper, we propose a novel graph clustering network to make full use of graph structural information. To capture the higher-order structural information, we design a graph mutual infomax module, effectively maximizing mutual information between graph-level and node-level representations, and employ a trinary self-supervised module that includes modularity as a structural constraint. Our proposed model outperforms many state-of-the-art methods on various datasets, demonstrating its superiority.
@@ 1 Introduction
Clustering plays a pivotal role in data mining, facilitating the categorization of real-world objects into clusters based on attributes or structural information without manual pre-labeling. Its practical applications are diverse. For instance, clustering users and products offers valuable insights for recommender platforms [11]. In biological networks [12], clustering contributes to unveiling intricate mechanisms in the human body. The application of clustering in social networks aids in the identification of terrorist organizations [15], contributing to the maintenance of regional security.
Clustering methods have undergone extensive evolution, striving to incorporate richer information. Initially, methods focus on either graph structure [7] or node attributes [20]. Later, researchers realize that achieving superior clustering performance requires the integration of node attributes and graph structure, which are crucial clues for capturing node similarities. As deep learning expands its applications across various domains [9], clustering methods can also leverage its ability to utilize information from both sides [17]. Notably, approaches based on graph convolutional networks (GCNs) [2, 16] aggregate neighbors’ features to obtain node representations and conduct clustering on the representations, resulting in more similar representations for nodes connected by shorter paths.
However, most existing methods are limited to handling higher-order structural information. For instance, in the Citeseer dataset [^1] Further details about the Citeseer dataset are provided in Section 5., 52.64% of nodes within the same cluster are connected through 7 or more hops. Merely employing GCN fails to capture this complex structural information. Thus, the challenge lies in how to capture the higher-order structural information of the graph. Additionally, some methods combine GCN with AE and employ a dual self-supervised module for training [1, 6]. These methods generate soft clustering assignments by measuring node distances to cluster centers derived from the encoder. These assignments undergo squaring and normalization to yield a highly confident target distribution, supervising the model’s training. However, the target distribution emphasizes node attributes, neglecting graph structure. Hence, a key challenge is how to utilize graph structure to supervise the target distribution.
In this paper, we propose a novel model that fully leverages the higher-order structural information of the graph and node attributes. The contributions of the paper are summarized as follows:
@@
  • Extensive experiments on real-world datasets demonstrate the superiority of our model compared with the state-of-the-art methods.
@@ 2 Related Work
Methods Jointly Modeling Graph Structure and Node Attributes. Recent research has recognized the importance of considering both graph structure and node attributes. DAEGC [18] employs a graph attentional autoencoder to capture neighbor features of nodes. SDCN [1] integrates GCN and autoencoder with a dual self-supervised module. Further, CaEGCN [6] proposes a cross-attention fusion module, merging GCN and autoencoder to enhance crucial information. These methods consider both structural information and node attributes, leading to improved clustering performance.

### 19 Mar 2024 v2

@@ Abstract
Clustering holds profound significance in data mining. In recent years, graph convolutional network (GCN) has emerged as a powerful tool for deep clustering, integrating both graph structural information and node attributes. However, most existing methods ignore the higher-order structural information of the graph. Evidently, nodes within the same cluster can establish distant connections. Besides, recent deep clustering methods usually apply a self-supervised module to monitor the training process of their model, focusing solely on node attributes without paying attention to graph structure. In this paper, we propose a novel graph clustering network to make full use of graph structural information. To capture the higher-order structural information, we design a graph mutual infomax module, effectively maximizing mutual information between graph-level and node-level representations, and design a trinary self-supervised module that includes modularity as a structural constraint. Our proposed model outperforms many state-of-the-art methods on various datasets, demonstrating its superiority.
@@ 1 Introduction
Clustering plays a pivotal role in data mining, facilitating the categorization of real-world objects into clusters based on attributes or structural information in an unsupervised way. Its practical applications are diverse. For instance, clustering users and products offers valuable insights for recommender platforms [11]. In biological networks [12], clustering contributes to unveiling intricate mechanisms in the human body. The application of clustering in social networks aids in the identification of terrorist organizations [15], contributing to the maintenance of regional security.
Clustering methods have evolved for a long time, aiming to incorporate richer information. Initially, methods focused on either graph structure [7] or node attributes [20]. However, researchers later realized that achieving superior clustering performance requires integrating both node attributes and graph structure, crucial for capturing node similarities. As deep learning expands its applications across various domains [9], clustering methods can leverage its ability to utilize information from both sides [17]. Notably, approaches based on graph convolutional networks (GCNs) [2, 16] aggregate neighbors’ features to obtain node representations and conduct clustering, resulting in more similar representations for nodes connected by shorter paths.
However, most existing methods are limited to handling higher-order structural information. For instance, in the Citeseer dataset, 52.64% of nodes within the same cluster are connected through 7 or more hops. Merely employing GCN fails to capture this complex structural information. Thus, the challenge lies in how to capture the higher-order structural information of the graph. Additionally, some methods combine GCN with autoencoder (AE) and employ a dual self-supervised module for training [1, 6]. These methods generate soft clustering assignments by measuring node distances to cluster centers derived from the encoder. However, the target distribution emphasizes node attributes, neglecting graph structure. Hence, a key challenge is how to utilize graph structure to supervise the target distribution.
In this paper, we propose a novel model that fully leverages the higher-order structural information of the graph and node attributes. Initially, an attribute-enriched GCN (AGCN) layer is constructed using a combination of encoder and GCN to aggregate neighbors’ information. To overcome the first challenge, a graph mutual infomax module is designed to capture the higher-order structural information of the graph. To overcome the second challenge, a trinary self-supervised module is proposed to optimize the entire model, taking both graph structure and node attributes into account. The contributions of the paper are summarized as follows:
@@ 2 Related Work
Methods Jointly Modeling Graph Structure and Node Attributes. Recent research has recognized the importance of considering both graph structure and node attributes. DAEGC [18] employs a graph attentional autoencoder to capture neighbor features of nodes. SDCN [1] integrates GCN and AE with a dual self-supervised module. Further, CaEGCN [6] proposes a cross-attention fusion module, merging GCN and autoencoder to enhance crucial information. These methods consider both structural information and node attributes, leading to improved clustering performance.